OpenStudy (anonymous):

Need some help with limits, why the limit of sen(f(x))/f(x) when (x,y)->(0,0) is 1???

OpenStudy (turingtest):

in English, do you mean why is\[\lim_{x\to0}\frac{\sin x}x\]?

OpenStudy (turingtest):

oh you have two variables, I see...

OpenStudy (anonymous):

yes, I have two variables I read that is always = 1 but I really don't get why :S

OpenStudy (turingtest):

it makes sense intuitively, right? since we know the above is true I'm just not sure how to prove that it's 1 in multivariable situations... is it not sufficient to say that\[\lim_{(x,y)\to(0,0)}\frac{\sin f(x)}{f(x)}\]and since \(y=f(x)\) we have\[\lim_{y\to0}\frac{\sin y}y\]on which we can use lhospital's rule? (probably not rigorous enough, sorry...)

OpenStudy (anonymous):

U_U what you write is true but I need to know how is this true in multivariable situations :S

OpenStudy (anonymous):

are you sure it aint? \[\lim _{(x,y)→(0,0)}sinf(x,y) \div f(x,y)\]

OpenStudy (anonymous):

yes you wrote it correctly sorry :S

OpenStudy (anonymous):

can you help me?

OpenStudy (anonymous):

well, lets think about it, are we allowed to evaluate the limits one after another?

OpenStudy (anonymous):

well if we have two variables yes, right?

OpenStudy (anonymous):

ok, just read the definition of limit, it must "evenly" converge along any of the possibilities so we could use lhospital in any of the cases

OpenStudy (anonymous):

so I use L'hopital twice?

OpenStudy (anonymous):

\[\lim _{(x,y)→(0,0)}sinf(x,y) \div f(x,y)\] , saly f(x,y)= x+y+1, it does not hold... since sen(1)/1 =not 1

OpenStudy (anonymous):

is the f(x,y) given?

OpenStudy (anonymous):

check the third post here: http://www.physicsforums.com/archive/index.php/t-318232.html

OpenStudy (anonymous):

ok I'll write the hole thing \[\lim_{(x,y) \rightarrow (0,0)}\sin (xy) \div xy\]

OpenStudy (anonymous):

ahhh ok, in that case, just change to polar coordinates, the lim (x,y)->0,0 turns into lim(r)->0

OpenStudy (anonymous):

Now I get it what you meant before about to use L'hopital in both cases I read that you can only use L'hopital in one variable but if I make that change I can use it thnx

OpenStudy (anonymous):

what I meant is that you evaluate lim x->0 (f(x,y_0)) as if y_0 where constant, and for every y_0 notequal 0 the lim must converge, and then the same but with y,x_0

OpenStudy (anonymous):

then again going to polar coordinates is the right thing to do here.

OpenStudy (anonymous):

This is what I did because x,y -> 0,0 I just make x=y so I'll just have one variable and then I just use L'hopital, Am I wrong?

OpenStudy (anonymous):

yes, it is wrong because you just tested one path, a function of 2 variables represents a surface, the lim must converge from every path to the point that its being tested

OpenStudy (anonymous):

when transforming to polar coordinates, you will get a function of r,theta,gamma, but since the limit simplifies to only r, you can use l'hospital o solve it

OpenStudy (anonymous):

mmm I see...

OpenStudy (anonymous):

as a completely different argument it is also valid to state: f(x)= x and so is f(y)=y, are continuous functions around 0. therefore f(x,y) =x*y is also continuous around 0,0, this implies that if one of the paths converges, all the others must also converge.

OpenStudy (anonymous):

thnx really now I get it